OpenStudy (anonymous):

f(t)=t^2-6t .. find the domain and sketch the graph.. can somebody explain me how to do this

OpenStudy (anonymous):

\[f(t)= t^2-6t\]

OpenStudy (anonymous):

When graphing parabolas, it is very useful to put the function into vertex form first then using transformations to determine where the graph is. Do you know about transformations of graphs?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

okay that's cool. If you want the simple way just use a graphing calculator. if not I will try to explain the transformation stuff

OpenStudy (anonymous):

i dont have a graphing calculator

OpenStudy (anonymous):

or you can find the vertex of the parabola and plot some points and guesstimate the graph

OpenStudy (anonymous):

okay I will try to teach you

OpenStudy (anonymous):

that is the graph of f(x)=x^2 other quadratic functions or transformations of that graph. for example, it's moved to the right/left, up/down, stretched...etc.

OpenStudy (anonymous):

when you put a quadratic in vertex form you are putting that function in the form f(x)=(x-h)+k where (h,k) is the vertex

OpenStudy (anonymous):

do you know about completing the square to solve quadratics?

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

so completing the square to solve x and doing it to put the function into vertex form is slightly different. More time sake, I will just show putting it into vertex form. so \[f(x)= t^{2}-6t\]

OpenStudy (anonymous):

you want to change this so that you will be able to factor it. To factor it however, you need a constant (the number after -6t) that is (half of 6) squared